OpenStudy (anonymous):

Let xy=1 and let dy/dt=5 Find dx/dt when x=5

OpenStudy (raspberryjam):

Okay so the first step is to take the derivative of xy=1. Don't forget the product chain rule! \[ xy=1\] \[ x\frac{dy}{dt}+y\frac{dx}{dt}=0\] Then just plug everything in! \[ (5)(5)+5\frac{dx}{dt}=0\] and solve for \[\frac{dx}{dt}\]

OpenStudy (anonymous):

then the answer would be -5. If i multiply the two five's together , i get 25. move it to the other side and divide by 5 i get -5. When i submit that answer on my end, it comes out wrong.

OpenStudy (anonymous):

The method was correct.

OpenStudy (raspberryjam):

Hmm...it should be right. Double check the numbers in the question?

OpenStudy (anonymous):

YEAH its still saying its wrong.

OpenStudy (raspberryjam):

Ahh wait my mistake. I accidentally put in 5 for y. So solve for y first. and put that in instead of the 5*dx/dt.

OpenStudy (anonymous):

\[y=\frac{1}{5}|x=5\]We're given the value for \(y\). As I said, the method is correct.